Please can you help with the following question?
Prove that cos(Ø+90º)=-sinØ

Answer by jim_thompson5910(35256)

cos(Ø+90)=-sin(Ø)

cos(Ø)cos(90)-sin(Ø)sin(90)=-sin(Ø)

cos(Ø)(0)-sin(Ø)(1)=-sin(Ø)

-sin(Ø)=-sin(Ø)


So the identity has been verified.